# CC-Diff: Enhancing Contextual Coherence in Remote Sensing Image Synthesis

### v2

@@ Abstract
Accurately depicting real-world landscapes in remote sensing (RS) images requires precise alignment between objects and their environment. However, most existing synthesis methods for natural images prioritize foreground control, often reducing the background to plain textures. This neglects the interaction between foreground and background, which can lead to incoherence in RS scenarios. In this paper, we introduce CC-Diff, a Diffusion Model-based approach for RS image generation with enhanced Context Coherence. To capture spatial interdependence, we propose a sequential pipeline where background generation is conditioned on synthesized foreground instances. Distinct learnable queries are also employed to model both the complex background texture and its semantic relation to the foreground. Extensive experiments demonstrate that CC-Diff outperforms state-of-the-art methods in visual fidelity, semantic accuracy, and positional precision, excelling in both RS and natural image domains. CC-Diff also shows strong trainability, improving detection accuracy by 2.04 mAP on DOTA and 2.25 mAP on the COCO benchmark. Code is available at https://github.com/AZZMM/CC-Diff.
@@ 1 Introduction
Recent studies have demonstrated the effectiveness of synthetic images in improving the performance of visual perception methods [37, 43, 56, 2]. This promise has motivated growing interest within the Geoscience community to explore remote sensing (RS) image synthesis [49, 12, 39, 36, 53], which has the potential to enhance the accuracy on various analytical tasks, e.g., oriented object detection [40, 15, 51] and landscape segmentation [21, 41, 3].
Despite remarkable advancements in visual quality, existing RS image generation methods [49, 12, 36] typically rely on text prompts to convey image semantics, which struggle to effectively describe the quantity, position, and orientation of foreground objects. Some studies tackle this issue by incorporating dense guidance information (e.g., semantic maps) to improve controllability [39, 53], but this inevitably increases annotation costs and limits the flexibility and diversity of the generated results.
Motivated by recent advances in controlled image generation, the Layout-to-Image (L2I) synthesis technique offers a promising solution to the aforementioned problem. However, existing L2I methods [57, 16, 59, 58, 47] primarily focus on ensuring the fidelity of foreground attributes concerning conditional information (e.g., texture, color, position) while overlooking their contextual alignment with the background (see Figure 1). This oversight may hold less significance in natural image generation, where the background primarily serves as a backdrop (or canvas) for object placement, allowing for minimal interaction between the two. In contrast, this lack of attention can be detrimental in RS images, where objects are closely tied to the semantics of their environment, making contextual coherence essential for generating semantically plausible representations (see Figure 1 for illustration).
To solve this problem, this study introduces CC-Diff, a Diffusion Model-based image generation method with enhanced Context Coherence. A closer examination of the internal design of current L2I models reveals that the issue likely stems from separate, non-interacting modules for foreground and background synthesis, and this lack of integration results in semantic inconsistencies in the generated images. While CC-Diff uses a multi-branch structure similar to existing methods for synthesizing distinct instances, it innovates by introducing a sequential process where background generation is conditioned on the features of the synthesized foreground elements. To further enhance the realism and coherence of the generated images, CC-Diff employs two sets of learnable queries to explicitly capture both the textural details of the background and its semantic relationship with the foreground instances. Extensive experiments on RS and natural images demonstrate that CC-Diff generates realistic images with high controllability and contextual coherence. Additionally, CC-Diff improves detection accuracy by 2.04 mAP on the popular DOTA dataset [46] in RS studies, and 2.25 mAP on the more generalized COCO benchmark [19], highlighting its strong trainability.
@@
  • We introduce CC-Diff, an L2I framework designed to improve contextual coherence in generated images, a problem we find crucial for RS images but largely overlooked by most existing methods.
  • A novel sequential pipeline is used to model the interdependence between instances and their surroundings, with distinct queries employed to capture both background texture and its relationship with the foreground.
  • Experiments demonstrate CC-Diff’s ability to generate realistic and coherent images across both RS and natural domains, which are shown to be effective augmentations for object detection tasks.
Figure 1: Illustration of contextual incoherencies in RS images synthesized by MIGC [59]. Layouts are shown in the top-right corner, object classes are labeled below, and incoherencies are highlighted with dashed yellow boxes for each figure.
@@ 2 Related Work
Controllable Image Generation. There are two main categories of controllable image generation methods: Text-to-Image (T2I) and Layout-to-Image (L2I), both focused on generating images that correspond to specific input conditions. T2I methods aim to produce images that reflect the semantics of the provided textual descriptions. During the prominence of methods based on Generative Adversarial Network (GAN) [30, 48, 55], DALL-E [28] was among the first to showcase the effectiveness of autoregressive frameworks, with subsequent studies [8, 52, 4] further enhancing fidelity and scalability. A different line of research employs the Diffusion Model [38, 11] alongside text prompts to achieve realistic T2I synthesis [29, 34, 23, 32].
In contrast to T2I, L2I approaches focus on generating images from instance layout to achieve spatial accuracy.With the rise of Diffusion Models, L2I research [47, 42, 57, 16, 50] has increasingly focused on integrating layout conditions into image generation. Recently, Multi-Instance Generation (MIG) in L2I has attracted growing interest [59, 58, 6, 45], with a focus on enhancing foreground instance attributes by separating the generation of foreground elements from background content.
Figure 2: An overview of CC-Diff’s pipeline. The central component of CC-Diff is the Dual Re-sampler, which leverages learnable queries to capture detailed features of the foreground ($\bm{q^{fg}}$), background ($\bm{q^{bg}}$), and notably their interdependencies ($\bm{q^{rel}}$). These queries are then passed to the Conditional Generation Module, where the background is generated based on the integration of synthesized foreground features ($\bm{R^{fused}}$), ensuring foreground-awareness throughout the process. Please zoom in for better details.
Remote Sensing Image Synthesis. While T2I synthesis for natural images has advanced significantly, research in the RS domain is still in its early stages. Among the initial efforts, Txt2Img-MHN [49] bridges the semantic gap between text prompts and RS images through hierarchical prototype learning. DiffusionSat [12] leverages LDM [33] to generate realistic RS images, controlling semantics through the embeddings of textual descriptions and numeric metadata. RSDiff [36] proposes to introduce a super-resolution network to further enhance image quality. Despite producing visually plausible RS images, these methods struggle to accurately control attributes such as the spatial layout of foreground instances. To address this, recent studies [53, 9, 35, 39] have incorporated additional guiding information, but none focus on ensuring the coherence between foreground and background.
@@ 3 Preliminaries
3.1 Latent Diffusion Model (LDM)
LDM [32] enhances the computational efficiency of vanilla Diffusion Models [11, 23, 29] by performing denoising in the latent space. After obtaining the latent representation $\mathbf{z}\in Z$ of the image $\mathbf{x}$, LDM seeks to learn a denoising autoencoder $\epsilon_{\bm{\theta}}$, which progressively generating less noisy data $\mathbf{z_{T-1}},\mathbf{z_{T-2}},\dots,\mathbf{z_{0}}$ from an initial sampled noise $\mathbf{z_{T}}$. In practice, $\epsilon_{\bm{\theta}}$ is trained to predict the step-wise noise $\bm{\epsilon}$ of the forward process, with the objective expressed as
$\min_{\bm{\theta}}\mathcal{L}_{LDM}=\mathbb{E}_{\mathbf{z},\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}[\|\bm{\epsilon}-\epsilon_{\bm{\theta}}(\mathbf{z}_{t},\mathbf{c},t)\|_{2}^{2}]$ (1)
where $t$ is a time step sampled from interval ${1,...,T}$ and $\mathbf{c}$ represents the prompt embedding, enabling semantic controllability via the cross-attention mechanism.
@@ 3.2 Problem Definition
let $\mathcal{L}=(\mathcal{P},\{c_{i},b_{i}\}_{i=1}^{N})$ represent the layout containing $N$ foreground objects, where the $i$-th object is annotated by the class label $c_{i}$ and positioned by the oriented bounding box $b_{i}=[x_{i},y_{i},w_{i},h_{i},{\alpha}_{i}]$ (i.e., coordinates of the top-left corner $(x_{i},y_{i})$, width $w_{i}$, height $h_{i}$, and orientation angle ${\alpha}_{i}$ (definition detailed in Supp. Mat. Sec 7). $\mathcal{P}$ serves as an auxiliary textual description of the image’s global semantics (please refer to Supp. Mat. Sec 8 for the detailed construction protocol). In this context, CC-Diff aims to learn a mapping $\mathcal{G}:(\mathcal{L},\mathbf{z})\rightarrow\mathbf{I}$, where $\mathbf{I}$ represents the generated image semantically aligned with $\mathcal{L}$, and $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the Gaussian noise.
@@ 4.1 Overview
The diagram of the proposed CC-Diff is illustrated in Figure 2. Specifically, CC-Diff employs learnable queries in the Dual Re-sampler to explicitly extract semantic information and relationships of key concepts, which are then passed to the Conditional Generation Module to guide the synthesis of both foreground instances and the background content. To conceptually implement the sequential pipeline described in Sec. 1, relating queries are utilized to capture the inherent connection between instances and their environment, while feature maps from synthesized foreground objects are fused into the background generation branch to produce an instance-aware background texture. Detailed introduction will be provided in the following subsections.
4.2 Associating BG with FG via Relating Query
As highlighted in Sec. 1, the primary distinction between L2I for natural and RS images is the level of textural complexity in the background and the critical role of contextual alignment with foreground instances. Accordingly, unlike most existing L2I approaches that use learnable queries solely to extract features of foreground objects (denoted as $\bm{q^{fg}}$), we introduce specialized queries to model background texture (background query $\bm{q^{bg}}$) and to capture the relationship between foreground and background (relating query $\bm{q^{rel}}$). The queries are learned through the proposed Dual Re-sampler, which will be detailed in the following.
Reference Image Retrieval. We first briefly describe how reference images are retrieved from the training set. Instances for each foreground category are cropped from training images with SAM[41] to build an exemplar dictionary, from which we randomly sample at generation time as image prompts. For background guidance, we sample one of the top five training images with the highest semantic similarity, computed using CLIP embeddings, as the background reference. Their embeddings are denoted as $\bm{f_{ins}}$ and $\bm{f_{img}}$, and please refer to Supp. Mat. Sec 9 for detailed retrieving approach.
Dual Re-sampler. As illustrated in Figure 2, the Dual Re-sampler comprises two distinct re-sampler modules for foreground and background (denoted as ‘FG-Resampler’ and ‘BG-Resampler’, respectively), along with an intermediate relating network that models the connection between the two. Both two Re-samplers consist of four consecutive combinations of Cross-Attention (CA) and Feed Forward Network (FFN) layers, where CA enables the interaction between input queries ($\bm{q^{fg}}$ and $\bm{q^{bg}}$) and the embedding of reference images ($\bm{f_{ins}}$ and $\bm{f_{img}}$) retrieved from the training set (further details will be provided later).
Specifically, the token of the $i$-th foreground instance (denoted as $\bm{h_{i}}$) by the FG-Resampler as follows
$\bm{h_{i}}=\operatorname{FG-Resampler}\left(\bm{q^{fg}},K_{i},V_{i}\right)$ (2)
where $K_{i},V_{i}$ $(i=1,...,N)$ are projections of $N$ reference instance embeddings $\bm{f_{ins}}$. In this way, the FG-Resampler serves as an image adapter, querying and extracting relevant features from reference instances to aid in the generation of realistic foreground objects.
To establish the connection between $\bm{q^{fg}}$ and $\bm{q^{bg}}$, a set of relating queries ($\bm{q^{rel}}$) is introduced to capture their contextual information, which is then integrated with $\bm{q^{bg}}$ to generate semantically coherent background textures. Specifically, $\bm{q^{rel}}$ interacts with the grounded foreground tokens through a Gated Self-Attention (GSA) [16] network, which can be formulated as
${\bm{q^{rel}}}={\bm{q^{rel}}}+\tanh(\gamma)\cdot\left(\operatorname{SelfAttn}\left(\left[\bm{{q^{rel}},{\bm{h^{e}}}}\right]\right)\right)$ (3)
where $\gamma$ is a learnable scalar and $\bm{h^{e}}=\left[{h^{e}_{1},...,h^{e}_{N}}\right]$ are grounded foreground tokens obtained through adding the bounding box formation into the output of FG Re-sampler. Subsequently, only the relating query components of the GSA output encoding the contextual semantics are retained, concatenated with the background query $\bm{q^{bg}}$, and then fed into the BG-Resampler, which can be expressed as
$\bm{h^{bg}}=\operatorname{BG-Resampler}\left(\left[\bm{q^{rel}},\bm{q^{bg}}\right],K^{bg},V^{bg}\right)$ (4)
where $K^{bg},V^{bg}$ are obtained from the embedding of reference image $\bm{f_{img}}$. Similar to the FG-Resampler, the BG-Resampler trains $\bm{q^{bg}}$ to extract features from the reference image, while considering the grounded context of the synthesized foreground objects, thereby enhancing the overall coherence of the generated image.
4.3 Foreground & Background Generation

### v3

@@ Abstract
Existing image synthesis methods for natural scenes focus primarily on foreground control, often reducing the background to simplistic textures. Consequently, these approaches tend to overlook the intrinsic correlation between foreground and background, which may lead to incoherent and unrealistic synthesis results in remote sensing (RS) scenarios. In this paper, we introduce CC-Diff, a Diffusion Model-based approach for RS image generation with enhanced Context Coherence. Specifically, we propose a novel Dual Re-sampler for feature extraction, with a built-in ‘Context Bridge’ to explicitly capture the intricate interdependency between foreground and background. Moreover, we reinforce their connection by employing a foreground-aware attention mechanism during the generation of background features, thereby enhancing the plausibility of the synthesized context. Extensive experiments show that CC-Diff outperforms state-of-the-art methods across critical quality metrics, excelling in the RS domain and effectively generalizing to natural images. Remarkably, CC-Diff also shows high trainability, boosting detection accuracy by 1.83 mAP on DOTA and 2.25 mAP on the COCO benchmark.
Figure 1: Comparison of (a) parallel and (b) conditional generation pipeline. The conditioning mechanism (denoted by red arrows) enhances the contextual conherence of generation results.
@@ 1 Introduction
Recent studies have highlighted the remarkable potential of synthetic imagery in boosting visual perception tasks [38, 46, 59, 3]. Motivated by these promising outcomes, the Geoscience community has devoted increasing attention to controllable remote sensing (RS) image synthesis [52, 14, 40, 37, 56], aiming to improve accuracy in a variety of analytical tasks [42, 16, 54, 22, 43, 4].
While many existing efforts [52, 14, 37] rely on textual prompts to encode image semantics, these prompts often fail to capture essential spatial cues (e.g., position and orientation) of foreground objects. To address this gap, some researchers incorporate dense guidance (e.g., semantic maps) to enhance controllability [40, 56], which inevitably raises annotation requirements and restricts both flexibility and diversity in the generated outputs.
Motivated by recent advances in spatially controllable image generation, the Layout-to-Image (L2I) technique offers a promising solution to the aforementioned challenge. However, most existing L2I methods [60, 17, 62, 61, 50] adopt a parallel generation pipeline (Figure 1 (a)), which primarily focuses on aligning the attributes of synthesized foreground instances (e.g., texture, color, position) with the given prompt, often overlooking broader contextual coherence with the background. This oversight may be less critical in many natural image scenarios as the background usually acts as a simple backdrop for object placement, it becomes far more consequential in RS imagery, where foreground objects are closely interlinked with their environment. Hence, ensuring contextual coherence is vital for producing semantically consistent results (see Figure 2).
To address this issue, we introduce CC-Diff, a Context-Coherent Diffusion Model for RS image generation. An in-depth analysis of existing L2I methods reveals that semantic inconsistencies primarily stem from separate, non-interacting modules handling foreground and background synthesis. While CC-Diff maintains a multi-branch design, it differs from existing approaches by establishing a cross-module conditioning mechanism that links the two synthesis pipelines. This mechanism enhances foreground awareness throughout both background feature extraction and rendering, as shown in Figure 1(b). Comprehensive experiments on RS datasets demonstrate that CC-Diff yields realistic, contextually coherent results with a high degree of controllability. In addition, CC-Diff exhibits strong trainability: it boosts detection accuracy by 1.83 mAP on the challenging DOTA dataset [49] and by 2.25 mAP on the more generalized COCO benchmark [20], underscoring its robust generalizability.
Figure 2: Illustration of contextual incoherencies in RS images synthesized by [62]. Layouts are shown in the top-right corner, object classes are labeled below with quotation marks, and incoherencies are highlighted with dashed yellow boxes.
@@
  • We introduce CC-Diff, an L2I framework originally conceived to address the incoherence in RS imagery synthesis, yet readily applicable to broader image domains.
  • We propose a conditional pipeline that explicitly captures the interdependence between foreground instances and their backgrounds, utilizing distinct learnable queries to integrate both background texture and the semantic context of the foreground.
  • Extensive experiments show that CC-Diff not only generates realistic, semantically consistent images across both RS and natural domains, but also serves as a highly effective augmentation strategy for object detection tasks.
@@ 2 Related Work
Controllable Image Generation. Text-to-Image (T2I) and Layout-to-Image (L2I) are the two main categories of controllable image generation methods. T2I approaches aim to synthesize images reflecting the semantics of textual descriptions. Early work in this field was dominated by Generative Adversarial Networks [32, 51, 58], until DALL-E [30] demonstrated the potential of autoregressive frameworks. Subsequent research [9, 55, 5] further improved fidelity and scalability. Meanwhile, another line of work employs Diffusion Models [39, 12] guided by text prompts to achieve realistic T2I synthesis [31, 35, 25, 33]. In contrast, L2I methods generate images from instance layouts to ensure spatial precision. With the rise of Diffusion Models, L2I research [50, 45, 60, 17, 53] increasingly focuses on integrating layout conditions. Recently, Multi-Instance Generation (MIG) [62, 61, 7, 48] has gained traction by separating the generation of foreground elements from background content, thereby enhancing their attributes.
Remote Sensing Image Synthesis. Although T2I for natural images has made significant progress, T2I in the RS domain remains in its infancy. Early attempts include Txt2Img-MHN [52], which employs hierarchical prototype learning to bridge the semantic gap between text prompts and RS imagery, and DiffusionSat [14], which leverages LDM [34] alongside textual and numeric metadata. RSDiff [37] further enhances image quality through super-resolution. While these methods produce visually plausible RS images, they struggle to precisely control attributes such as the spatial layout of foreground instances. The contemporary study AeroGen [41] introduces a layout-controllable diffusion model supporting rotated bounding boxes, and related efforts [56, 10, 36, 40] add additional guiding information. However, none explicitly address maintaining coherence between foreground and background.
@@ 3 Preliminaries
3.1 Latent Diffusion Model
The Latent Difffusion Model (LDM) [33] enhances the computational efficiency of vanilla Diffusion Models [12, 25, 31] by performing denoising in the latent space, and improving controllability through the integration of a cross-attention mechanism. After obtaining the latent representation $\bm{z}\in Z$ of the image $\bm{x}$, LDM seeks to learn a denoising autoencoder $\epsilon_{\bm{\theta}}$, which progressively generating less noisy data $\bm{z_{T-1}},\bm{z_{T-2}},\dots,\bm{z_{0}}$ from an initial sampled noise $\bm{z_{T}}$. In practice, $\epsilon_{\bm{\theta}}$ is trained to predict the step-wise noise $\bm{\epsilon}$ of the forward process, with the objective expressed as
$\min_{\bm{\theta}}\mathcal{L}_{LDM}=\mathbb{E}_{\bm{z},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),t}[\|\bm{\epsilon}-\epsilon_{\bm{\theta}}(\bm{z}_{t},\bm{c},t)\|_{2}^{2}]$ (1)
where $t$ is a time step sampled from interval ${1,...,T}$ and $\bm{c}$ represents the prompt embedding.
3.2 Cross-Attention Mechanism
In LDM, the semantics of generated images are guided by conditional prompts through a cross-attention (CA) mechanism [33]. Given a latent image feature $\mathbf{f}$ and a condition $\mathbf{c}$, they are first projected to obtain the query, key, and value representations: $\mathbf{Q}=\mathbf{W_{Q}}\cdot\mathbf{f}$, $\mathbf{K}=\mathbf{W_{K}}\cdot\mathbf{c}$, and $\mathbf{V}=\mathbf{W_{V}}\cdot\mathbf{c}$, where $\mathbf{W_{Q}}$, $\mathbf{W_{K}}$, and $\mathbf{W_{V}}$ are learnable projection matrices. The CA between $\mathbf{f}$ and $\mathbf{c}$ is then computed by
$\text{CA}(\mathbf{f},\mathbf{c})=\text{Softmax}\left(\frac{\mathbf{Q}(\mathbf{f})\mathbf{K}(\mathbf{c})^{\top}}{\sqrt{d}}\right)\mathbf{V}(\mathbf{c}),$ (2)
All subsequent cross-attention operations in this paper follow the formulation in Eq. 3.
@@ 3.3 Problem Definition
let =$\mathcal{L}=\{\mathbf{c_{i}},\mathbf{b_{i}}\}_{i=1}^{N}$ represent the layout containing $N$ foreground objects, where the $i$-th object is assigned a class label $\mathbf{c_{i}}$ and an oriented bounding box $\mathbf{b_{i}}=[x_{i},y_{i},w_{i},h_{i},{\theta}_{i}]$. Specifically, $(x_{i},y_{i})$ denotes the coordinates of the top-left corner, $w_{i}$ and $h_{i}$ are the width and height, and ${\theta}_{i}$ is the orientation angle. Moreover, an auxiliary textual description $\mathcal{P}$ is introduced to capture the image’s global semantics. In this setting, CC-Diff learns a mapping $\mathcal{G}:(\mathcal{L},\mathcal{P},\mathbf{z})\rightarrow\bm{I}$, where $\bm{I}$ represents the generated image semantically aligned with $\mathcal{L}$ and $\mathcal{P}$, and $\mathbf{z}$ denotes the Gaussian noise. For details of the construction of $\mathcal{L}$ and $\mathcal{P}$, please refer to Sec. 7 and Sec. 8 of the Appendix.
Figure 3: The framework of CC-Diff. The Dual Re-sampler extracts condition embeddings from user inputs (bounding boxes, class labels, and descriptions), guiding the Conditional Generation Module (CGM) to produce contextually coherent outputs.
Figure 4: The architecture of Dual Re-sampler. The context query $\mathbf{q^{ctx}}$ obtains contextual semantics of FG objects from the FG Re-sampler, then incorporates them into BG feature extraction within the BG Re-sampler, thereby establishing the FG-BG association.
Figure 5: The architecture of Conditional Generation Module (CGM). The BG feature is rendered using the fused FG representation $\mathbf{R^{fused}}$, ensuring FG-awareness throughout generation.
@@ 4.1 Overview
As illustrated in Figure 3, CC-Diff employs a Dual Re-sampler to encode the semantic information of foreground (FG) and background (BG), along with their relationships, into condition embeddings. These embeddings are then passed to the Conditional Generation Modules (CGMs) within the UNet of LDM to guide the generation with enhanced contextual coherence. Detailed introductions of these two modules are provided as follows.
4.2 Dual Re-sampler for FG-BG Association
As its name suggests, the Dual Re-sampler comprises two specialized re-sampler modules for extracting FG and BG features (denoted as ‘FG Re-sampler’ and ‘BG Re-sampler’, respectively). To model their underlying dependencies, a relating component named Context Bridge is involved to explicitly establish the connection between the two re-samplers. An overview of the Dual Re-sampler architecture is provided in Figure 4.
FG Re-sampler. Inspired by the Perceiver architecture [1], the FG Re-sampler utilizes a set of learnable queries $\mathbf{q^{fg}}\in\mathbb{R}^{1\times N_{q}\times d}$ (where $N_{q}$ denotes the number of query tokens and $d$ the latent dimension) to extract the semantic information of foreground objects. Specifically, given the text embedding of object labels $\mathbf{f^{obj}}\in\mathbb{R}^{N\times S\times d}$ (computed by a pre-trained CLIP encoder [29] and $S$ denotes the sequence length), the FG Re-sampler comprises $L$ attention layers, each consisting of a Cross Attention (CA) block followed by a Feed Forward Network (FFN). Formally, the CA mechanism in FG Re-sampler is expressed as:
$\text{CA}(\mathbf{q^{fg}},\mathbf{f^{obj}})=\text{Softmax}\left(\frac{\mathbf{Q}(\mathbf{\mathbf{q^{fg}}})\mathbf{K}(\mathbf{f^{obj}})^{\top}}{\sqrt{d}}\right)\mathbf{V}(\mathbf{f^{obj}}),$ (3)
where the output of each layer provides the query input for the subsequent layer, ultimately yielding the FG token $\mathbf{h^{fg}}\in\mathbb{R}^{N\times N_{q}\times d}$.
Beyond capturing semantic information, the grounding token $\mathbf{h^{bbox}}\in\mathbb{R}^{N\times 1\times d}$ encodes the spatial distribution of objects by mapping bounding box parameters through a Fourier Embedding module [24]. The element-wise sum of $\mathbf{h^{fg}}$ and $\mathbf{h^{bbox}}$ then serves as the comprehensive FG representation passed on to the subsequent processing stages.
Context Bridge (CB). To bridge FG and BG and improve coherence, we additionally introduce a learnable context query $\mathbf{q^{ctx}}\in\mathbb{R}^{1\times N_{q}\times d}$. Specifically, $\mathbf{q^{ctx}}$ is fed into a Self-Attention (SA) block together with the previously learned FG tokens (i.e., $\mathbf{h^{fg}}$ and $\mathbf{h^{bbox}}$) to capture cross-token relationships and contextual cues. The resulting output is then added back to $\mathbf{q^{ctx}}$, yielding the final output of the Context Bridge, denoted as $\text{CB}(\mathbf{q^{ctx}})$. Mathematically, $\text{CB}(\mathbf{q^{ctx}})$ can be expressed as
$\text{CB}(\mathbf{q^{ctx}})=\mathbf{q^{ctx}}+\tanh(\gamma)\cdot\text{TS}(\text{SA}([\mathbf{q^{ctx}},\mathbf{h^{fg}}+\mathbf{h^{bbox}}])),$ (4)
where $\text{TS}(\cdot)$ is a Token Selection operation [17] that retains only the context query from the SA output, []delimited-[][$\cdot$][ ⋅ ] represents the concatenating operation and $\gamma$ is a learnable scalar. By extracting the contextual cues of FG instances, $\mathbf{q^{ctx}}$ further enriches the context-aware processing of BG features.
BG Re-sampler. Following the design of the FG Re-sampler, the BG Re-sampler also employs a learnable query $\mathbf{q^{bg}}\in\mathbb{R}^{1\times N_{q}\times d}$ to extract global semantics information from the textural description $\mathcal{P}$ for background synthesis. However, unlike the FG Re-sampler, it additionally incorporates grounded context from FG objects through $\text{CB}({\mathbf{q^{ctx}}})$. Concretely, the concatenated query ($[\mathbf{q^{bg}},\text{CB}({\mathbf{q^{ctx}}})]$) engages in the CA mechanism with the text embedding of $\mathcal{P}$ (denoted as $\mathbf{f^{all}}\in\mathbb{R}^{1\times S\times d}$). As a result, the output BG token $\mathbf{h^{bg}}\in\mathbb{R}^{1\times 2N_{q}\times d}$ encodes BG semantics with the awareness of FG context.
Condition Embeddings. The final outputs of the Dual Re-sampler are the condition embeddings for the $N$ FG objects ($\{\mathbf{e^{fg}_{i}}\}_{i=1}^{N}$) and the BG ($\mathbf{e^{bg}}$). These embeddings are constructed by concatenating the corresponding learned tokens, as follows:
$\mathbf{e^{fg}_{i}}$ $=[\mathbf{h_{i}^{fg}},\mathbf{h_{i}^{bbox}},\texttt{cls}(\mathbf{c_{i}})]\in\mathbb{R}^{1\times(N_{q}+2)\times d}$ (5)
$\mathbf{e^{bg}}$ $=[\mathbf{h^{bg}},\texttt{cls}(\mathcal{P})],$ (6)
where $\mathbf{h_{i}^{fg}}$, $\mathbf{h_{i}^{bbox}}$ are the token slices pertaining to the $i$-th object, and $\texttt{cls}(\cdot)$ denotes the output cls token produced by the CLIP text encoder.